# DF-Calib: Targetless LiDAR-Camera Calibration via Depth Flow

### v1

@@ Abstract
Precise LiDAR-camera calibration is crucial for integrating these two sensors into robotic systems to achieve robust perception. In applications like autonomous driving, online targetless calibration enables a prompt sensor misalignment correction from mechanical vibrations without extra targets. However, existing methods exhibit limitations in effectively extracting consistent features from LiDAR and camera data and fail to prioritize salient regions, compromising cross-modal alignment robustness. To address these issues, we propose DF-Calib, a LiDAR-camera calibration method that reformulates calibration as an intra-modality depth flow estimation problem. DF-Calib estimates a dense depth map from the camera image and completes the sparse LiDAR projected depth map, using a shared feature encoder to extract consistent depth-to-depth features, effectively bridging the 2D-3D cross-modal gap. Additionally, we introduce a reliability map to prioritize valid pixels and propose a perceptually weighted sparse flow loss to enhance depth flow estimation. Experimental results across multiple datasets validate DF-Calib’s accuracy and generalization, with DF-Calib achieving a mean translation error of $0.635cm$ and rotation error of $0.045^{\circ}$ on the KITTI dataset.
I Introduction
@@
In modern robotic systems, such as autonomous driving, LiDAR and cameras are widely configured for their complementary capabilities. LiDAR provides high-precision 3D spatial measurement, while the camera captures detailed texture information. The fusion of the LiDAR point cloud and camera image is essential for advancing pivotal robotic technologies, e.g., SLAM [1, 2] and object detection [3, 4], ultimately improving perception accuracy and robustness. A fundamental prerequisite in this process is precise extrinsic calibration, which ensures accurate spatial alignment of heterogeneous sensor data. General calibration methods [5, 6] typically rely on handcrafted targets, such as checkerboards. However, in real-world scenarios, mechanical vibrations and operational disturbances cause gradual deviation in extrinsic parameters, necessitating frequent recalibration. The reliance on specific targets makes target-based calibration cumbersome and impractical in dynamic environments. Therefore, developing online targetless calibration methods is crucial for enhancing adaptability and usability in real-world applications.
However, compared to offline target-based methods, online targetless calibration relies on natural environmental features for registration due to the absence of predefined calibration targets. This exacerbates the challenge of bridging the significant modality gap between LiDAR point clouds and camera images, making it more difficult to achieve accurate extrinsic parameter estimation. Traditional online targetless methods [7, 8, 9] optimize extrinsic parameters by leveraging features such as edges, lane markings, and motion trajectories. However, their performance is constrained by the availability and consistency of these features, which may be sparse or unreliable in natural fields. As a result, these approaches become less effective in scenarios where preset cues are either insufficient or exhibit high variability.
To address this challenge, deep learning techniques have been introduced to reduce reliance on predefined targets and idealized conditions. Current learning-based methods [10, 11] typically extract features separately from the sparse depth map projected from the LiDAR point cloud and the camera image. Various approaches [12, 13] adopt an end-to-end framework, directly regressing extrinsic parameters with high accuracy but relying on specific camera configurations, limiting generalizability. In contrast, matching-based methods [14, 15] establish 2D-3D correspondences, decoupling extrinsic estimation from camera parameters and enhancing generalization. However, inherent modality differences lead to feature distribution discrepancies, making both paradigms struggle with consistent feature extraction and robust cross-modal alignment, ultimately compromising calibration accuracy. Furthermore, they fail to prioritize salient regions, such as edges or other structural areas, within the LiDAR and camera data, weakening critical feature influence and further degrading calibration accuracy. Recently, CalibDepth [16] reduces the modality gap by matching the sparse LiDAR-projected depth map with the camera estimated dense depth map. However, the LiDAR sparse depth map lacks the detailed geometric structures (e.g., edges and surfaces) that are present in the camera dense depth map, leading to registration challenges that ultimately compromise calibration accuracy.
In this paper, we propose DF-Calib, a matching-based targetless LiDAR-camera calibration method designed to address the aforementioned problems via depth flow. We employ monocular geometry estimation to generate a dense depth map from the camera image and complete the sparse LiDAR-projected depth map into a dense representation, leveraging a shared feature encoder to extract consistent depth-to-depth features. Additionally, we introduce a reliability map to guide the model prioritizing features extracted from valid pixels, thus minimizing the impact of noise introduced during depth completion. Subsequently, we estimate a depth flow that incorporates 2D-3D correspondences between the camera and LiDAR data, followed by a PnP solver to calculate the extrinsic parameter. To further improve the accuracy of the depth flow, we propose the perceptually weighted sparse flow (PWSF) loss, which encourages the model to focus on pixels within the salient regions of the depth map. Our main contributions are listed as follows:
@@
  • We propose DF-Calib, a targetless LiDAR-camera calibration method that transforms the calibration task into an intra-modality depth flow estimation problem, achieving high calibration accuracy while improving generalization.
  • We convert the camera image into a dense depth map and complete the sparse LiDAR depth map, employing a shared feature encoder to extract consistent depth-to-depth features, which effectively bridges the cross-modal gap and enhances calibration accuracy.
  • We introduce a reliability map to prioritize features from valid pixels and propose a perceptually weighted sparse flow loss to focus on salient regions of the depth maps, enhancing depth flow and calibration accuracy.
@@
  • Our experimental results across multiple datasets, including KITTI Odometry, KITTI Raw, and KITTI-360, demonstrate that DF-Calib achieves accurate extrinsic calibration results and exhibits reliable generalization ability.
II Related Work
@@
LiDAR-camera calibration can be broadly classified into target-based and targetless methods. While learning-based approaches are gaining prominence for their calibration accuracy and attracting considerable attention, we consider them a distinct category for further discussion.
II-A Target-based Methods
@@
Target-based LiDAR-camera calibration relies on calibration targets detectable in both modalities to establish accurate 2D-3D correspondences. Early methods [5, 6, 17] were inspired by camera intrinsic calibration using chessboard patterns. Over time, various calibration targets, such as plane boards with holes [18], triangles [19], polygons [20], spheres [21], and ArUco [22], have been developed to enhance feature extraction accuracy. While these methods improve calibration precision, the complex target designs increase production difficulty. Despite the high precision achievable, the reliance on targets introduces limitations related to accessibility and automation, making it unsuitable for online calibration.
II-B Targetless Methods
@@
Targetless methods eliminate the need for calibration targets by leveraging common features and generally establishing correspondences directly from natural scene structures. Frequently used features include edges [7] and lane markings [8], with extrinsic parameters optimized to maximize feature consistency.
@@
To avoid relying on predefined features, Nagy et al. [23] recover 3D geometry using Structure-from-Motion (SfM) and jointly optimize the extrinsic during the reconstruction process. Mutual information techniques [24, 25] align projected depth or intensity maps with camera images, optimizing the extrinsic by maximizing mutual information. Taylor et al. [9] employ motion-based approaches by matching trajectories but require diverse motion data for accurate optimization.
@@
In summary, targetless methods use a range of features and metrics to formulate an optimization problem for the extrinsic. However, their effectiveness is highly dependent on initial conditions, assumption validity, and scene suitability.
II-C Learning-based Methods
@@
Regression-based methods employ an end-to-end approach where the network directly regresses extrinsic parameters. RegNet [10] pioneers this framework, defining a pipeline of feature extraction, matching, and global regression. CalibNet [11] further introduces geometric constraints to improve consistency. CalibRCNN [26] and CALNet [27] refine these techniques by incorporating temporal losses for improved stability. CalibDepth [16] enhances feature matching between the sparse LiDAR-projected depth map and the estimated dense depth map but struggles to align complex geometric structures (e.g., lines and surfaces), reducing calibration accuracy. Despite their effectiveness, regression-based methods strongly depend on camera configurations, requiring retraining or fine-tuning for new sensor setups.
@@
Matching-based methods, in contrast, decouple feature matching from pose estimation, establishing 2D-3D correspondences and formulating calibration as a PnP problem. This improves generalization across different camera configurations. CFNet [14] introduces calibration flow, leveraging the EPnP algorithm [28], while DXQ-Net [15] integrates a differentiable pose estimation module with uncertainty filtering. However, both regression-based and matching-based approaches face challenges in handling cross-modal inconsistencies, particularly in extracting consistent features, ultimately restricting calibration accuracy. Additionally, they generally overlook salient regions in the LiDAR and camera data, diminishing the impact of critical features and further degrading calibration performance.
Unlike existing methods, our approach converts both camera images and LiDAR point clouds into dense depth maps, from which we extract consistent depth-to-depth features. By estimating the depth flow between these maps, we establish 2D-3D correspondences and solve for extrinsic parameters via a PnP solver, effectively addressing the cross-modal challenge. Additionally, we employ PWSF loss to emphasize salient regions in the depth map, enhancing depth flow accuracy and calibration precision.
Figure 1: Our proposed method first employs MOGE to estimate the dense depth map $D_{I}^{d}$ from the camera image. Simultaneously, the LiDAR point cloud is projected onto the image plane, generating a sparse depth map $D_{L}^{s}$, further densifying into $D_{L}^{d}$. These three depth maps are input into the network to estimate a depth flow encapsulating the 2D-3D correspondences between the camera and LiDAR data. Ultimately, a 6-DoF transformation $\mathbf{T}_{est}$ is computed via PnP with RANSAC.
III Proposed Method
@@
The inputs of the proposed method are a camera image and a LiDAR point cloud. The camera image is processed by MoGe [29] to estimate a dense depth map, while the LiDAR point cloud is projected into a sparse depth map and subsequently completed to a dense one. These three depth maps are fed into the network, which estimates a depth flow representing the 2D-3D correspondences between the camera and LiDAR data. Finally, the extrinsic parameter is estimated by the PnP solver. An overview of our method is shown in Fig. 1.
III-A Problem Formulation
@@
Given a misaligned camera image $I\in\mathbb{R}^{3\times W\times H}$ from the camera and LiDAR point cloud $P_{L}\in\mathbb{R}^{3\times N}$, the extrinsic calibration aim to estimate a 6-DoF extrinsic parameter $\mathbf{T}_{est}=\begin{bmatrix}\mathbf{R}_{est},\mathbf{t}_{est}\end{bmatrix}\in SE(3)$ that adjusts the initial extrinsic parameter $\mathbf{T}_{init}\in SE(3)$ for accurate sensor alignment. Here, $W$ and $H$ represent the image width and height, respectively. $\mathbf{R}_{est}\in SO(3)$ denotes the rotation matrix, and $\mathbf{t}_{est}\in\mathbb{R}^{3\times 1}$ represents the translation vector. For matching-based methods, the 2D-3D correspondences between image pixels and LiDAR points are first established. The extrinsic parameter is then estimated via the PnP solver, with RANSAC to prune outliers and enhance robustness.
III-B Depth Generation
The main challenge in LiDAR-camera extrinsic calibration lies in cross-modal data matching due to structure and information density differences between point clouds and images. To address this, we propose a novel method that transforms both modalities into dense depth maps, aligning their structure and information density for effective matching. First, we employ MoGe [29] to estimate the dense depth map $D_{I}^{d}\in\mathbb{R}^{1\times W\times H}$ from a camera image. Meanwhile, leveraging the camera’s intrinsic parameter $\mathbf{K}\in\mathbb{R}^{3\times 3}$ and initial extrinsic parameter $\mathbf{T}_{init}=\begin{bmatrix}\mathbf{R}_{init},\mathbf{t}_{init}\end{bmatrix}$, the LiDAR point cloud $P_{L}$ is projected onto the pixel coordinate system, resulting in a sparse depth map $D_{L}^{s}\in\mathbb{R}^{1\times W\times H}$. The projection process can be formulated as
$d_{i}\begin{bmatrix}u_{i}\\ v_{i}\\ 1\end{bmatrix}=\mathbf{KT}_{init}\begin{bmatrix}x_{i}\\ y_{i}\\ z_{i}\\ 1\end{bmatrix}=\mathbf{K}\begin{bmatrix}\mathbf{R}_{init}&\mathbf{t}_{init}\\ \mathbf{0}&1\end{bmatrix}\begin{bmatrix}x_{i}\\ y_{i}\\ z_{i}\\ 1\end{bmatrix},$ (1)
where $d_{i}$ represents the depth of each pixel, and $[x_{i},y_{i},z_{i},1]^{\mathrm{T}}$ is the homogeneous coordinate of the 3D point, while $[u_{i},v_{i},1]^{\mathrm{T}}$ is the corresponding pixel’s homogeneous coordinate in 2D. However, due to the inherent sparsity of LiDAR data, $D_{L}^{s}$ contains only a limited number of valid pixels, leading to a significant information density disparity between $D_{L}^{s}$ and $D_{I}^{d}$, which adversely affects subsequent depth flow estimation. To mitigate this issue, we employ an efficient depth completion method [30] to transform $D_{L}^{s}$ into a dense depth map $D_{L}^{d}\in\mathbb{R}^{1\times W\times H}$. As shown in Fig. 2, this process ensures that both the LiDAR point cloud and the camera image are effectively converted into dense depth maps, facilitating the subsequent depth flow estimation.
@@ III-B Depth Generation
Figure 2: Comparison of the sparse depth map projected by LiDAR, the dense depth map after depth completion, and the dense depth map estimated from the camera image. After completion, the depth maps generated by the LiDAR and the camera exhibit similar textures.

### v2

# UniCalib: Targetless LiDAR-Camera Calibration via Probabilistic Flow on Unified Depth Representations

@@ Abstract
Online targetless extrinsic LiDAR-camera calibration is essential for robust perception in computer vision applications such as autonomous driving. However, existing methods struggle with the significant modality gap between heterogeneous sensors and fail to handle unreliable correspondences arising from real-world challenges like occlusions and dynamic objects. To address these issues, we introduce UniCalib, a novel method that performs calibration by estimating a probabilistic flow on unified depth representations. UniCalib first bridges the modality gap by converting both the camera images and the sparse LiDAR points into unified, dense depth maps, enabling a unified encoder to learn consistent features. Subsequently, it learns a probabilistic flow field that captures the correspondence uncertainty to improve robustness. This probabilistic approach is reinforced by a reliability map and a perceptually weighted sparse flow loss, which guide the model to suppress the influence of unreliable regions. Experimental results on three datasets validate the accuracy and generalization of UniCalib. In particular, it achieves a mean translation error of $0.550\mathrm{cm}$ and a rotation error of $0.044^{\circ}$ on the KITTI dataset.
1 Introduction
Figure 1: Architectural comparison for LiDAR-camera calibration. (a) Modality-specific encoders followed by the fusion module. (b) Our method converts both modalities into dense depth maps, enabling a unified encoder.
@@ 1 Introduction
The fusion of camera and Light Detection And Ranging (LiDAR) data is essential for computer vision applications such as autonomous driving and 3D object detection [13, 8], owing to their complementary strengths in capturing texture and geometry. This requires precise extrinsic calibration to ensure accurate spatial alignment between the sensors. Traditional offline calibration methods [35, 27] rely on handcrafted targets, such as checkerboards. However, in real-world scenarios, mechanical vibrations and operational disturbances cause gradual deviation in the extrinsics, necessitating frequent recalibration. This dependence on specific physical targets is cumbersome and impractical in dynamic environments, making the development of online targetless calibration methods crucial for real-world applications.
Compared to target-based methods, online targetless calibration relies on structured environmental features for registration, requiring consistent and robust cross-modal feature matching between LiDAR and camera data. This is inherently challenging due to the heterogeneous nature of the two modalities. Traditional online targetless methods [1, 17, 25] bridge the modality gap by leveraging features such as edges, lane markings, and motion trajectories. However, their performance heavily depends on specific scene structures, making them less robust in complex or unstructured environments.
To overcome the reliance on predefined targets, deep learning techniques have been introduced. These methods generally follow two paradigms. End-to-end approaches [32, 18] directly regress extrinsics, but their network architectures are coupled with specific camera intrinsics, which limits their generalizability. In contrast, matching-based methods [16, 10] first establish 2D-3D correspondences, decoupling the learning process from camera parameters and thus enhancing generalization. However, as shown in Fig. 1 (a), a common limitation in both paradigms is the independent extraction of features from the camera image and the LiDAR-projected sparse depth map. Due to inherent modality differences, the extracted features exhibit a significant domain gap. Consequently, these methods struggle with robust cross-modal alignment, which ultimately compromises calibration accuracy. Furthermore, these methods often treat all potential correspondences with uniform importance. This inherent limitation makes them highly vulnerable to real-world challenges, such as dynamic objects and occlusions, as they lack a mechanism to distinguish reliable data from resulting outliers, ultimately degrading calibration accuracy.
In this paper, we propose UniCalib, a novel matching-based method that transforms the cross-modal calibration task into a probabilistic intra-modality depth flow estimation problem. As illustrated in Fig. 1 (b), we bridge the modality gap by converting both camera and LiDAR data into dense depth maps, thereby unifying them into the same modality. This enables a unified encoder to learn consistent features from the shared domain. To handle real-world uncertainty, our framework predicts a probabilistic depth flow field representing a per-pixel distribution over correspondences, guided by a reliability map that prioritizes valid LiDAR data. This probabilistic approach is reinforced by our novel Perceptually Weighted Sparse Flow (PWSF) loss, which utilizes the predicted uncertainty and an outlier probability map to suppress supervision in unreliable regions (dynamic objects and occlusions), thereby preventing overfitting adaptively. Finally, the robust 2D-3D correspondences derived from this process are fed into a PnP solver to compute the extrinsics.
Our main contributions are listed as follows:
@@
  • We propose UniCalib, a novel targetless LiDAR-camera calibration method that reframes the 2D-3D correspondence estimation task as a probabilistic intra-modality depth flow estimation problem, addressing the cross-modal challenge.
  • We convert both camera images and sparse LiDAR point clouds into the same modality, which enables a unified feature encoder to learn consistent representations, effectively bridging the modality gap.
  • We introduce a probabilistic depth flow network featuring a reliability map and a novel PWSF loss, which handles correspondence uncertainty and adaptively suppresses supervision in unreliable regions.
@@
  • The experimental results across multiple datasets, including KITTI Odometry, KITTI Raw, and KITTI-360, demonstrate that UniCalib achieves accurate extrinsic calibration results and exhibits reliable generalization ability.
Figure 2: Overview of UniCalib. Depth Generation converts camera and LiDAR inputs into dense depth maps, which are passed to the Probabilistic Calibration Network to estimate a probabilistic depth flow field. The mean of the predicted distribution, referred to as the estimated flow, is then used by the Extrinsic Estimation module to compute the LiDAR-camera extrinsics.
2 Related Work
@@
LiDAR-camera extrinsic calibration can be broadly classified into target-based and targetless methods. While learning-based approaches are gaining prominence for their calibration accuracy and attracting considerable attention, we consider them a distinct category for further discussion.
2.1 Target-based Methods
@@
Target-based LiDAR-camera calibration relies on calibration targets detectable in both modalities to establish accurate 2D-3D correspondences. Early methods [35, 27, 6] are inspired by camera intrinsic calibration using chessboard patterns. Over time, various calibration targets, such as plane boards with holes [33], triangles [3], polygons [14], spheres [34], and ArUco [4], have been developed to enhance feature extraction accuracy. While these methods improve calibration precision, the complex target designs increase production difficulty. Despite the high precision achievable, the reliance on targets introduces limitations related to accessibility and automation, making it unsuitable for online calibration.
2.2 Targetless Methods
@@
Targetless methods eliminate the need for calibration targets by leveraging special features and generally establishing correspondences directly from natural scene structures. Frequently used features include edges [1] and lane markings [17], with extrinsics optimized to maximize feature consistency.
@@
To avoid relying on predefined features, Nagy et al. [19] recover 3D geometry using Structure-from-Motion (SfM) and jointly optimize the extrinsics during the reconstruction process. Mutual information techniques [20, 7] align projected depth or intensity maps with camera images, optimizing the extrinsics by maximizing mutual information. Taylor and Nieto [25] employ motion-based approaches by matching trajectories but require diverse motion data for accurate optimization.
@@
In summary, targetless methods use a range of features and metrics to formulate an optimization problem for the extrinsics. However, their effectiveness is highly dependent on initial conditions, assumption validity, and scene suitability.
2.3 Learning-based Methods
@@
Regression-based methods employ an end-to-end approach where the network directly regresses extrinsics. RegNet [21] pioneers this framework, defining a pipeline of feature extraction, matching, and global regression. CalibNet [9] further introduces geometric constraints to improve consistency. CalibRCNN [23] and CALNet [22] refine these techniques by incorporating temporal losses for improved stability. CalibDepth [37] enhances feature matching between the sparse LiDAR-projected depth map and the estimated dense depth map but struggles to align complex geometric structures, reducing calibration accuracy. Despite their effectiveness, regression-based methods strongly depend on camera configurations, requiring retraining or fine-tuning for new sensor setups.
@@
Matching-based methods, in contrast, decouple feature matching from extrinsic estimation, establishing 2D-3D correspondences and formulating calibration as a PnP problem. This improves generalization across different camera configurations. CFNet [16] introduces calibration flow, leveraging the EPnP algorithm [12], while DXQ-Net [10] integrates a differentiable pose estimation module with uncertainty filtering. However, both regression-based and matching-based approaches face fundamental limitations. They struggle to extract consistent features across the modality gap and lack robust mechanisms to handle real-world outliers from sources like dynamic objects or occlusions, leaving them susceptible to spurious correspondences.
Diverging from conventional methods, our approach recasts the cross-modal problem into an intra-modality task by converting both camera images and LiDAR point clouds into dense depth maps. We then estimate a probabilistic depth flow between these maps, utilizing our innovative PWSF loss. This loss function incorporates predicted uncertainty and an outlier probability map to minimize the impact of unreliable areas. This process yields reliable 2D-3D correspondences, from which a PnP solver computes the final extrinsics.
3 Proposed Method
@@
An overview of our method is shown in Fig. 2. The inputs are a camera image and a LiDAR point cloud from the same scene. The camera image is processed by MoGe [28] to estimate a dense depth map, while the LiDAR point cloud is projected into a sparse depth map and subsequently completed to a dense one. These two dense depth maps are then fed into our probabilistic calibration network, with the sparse LiDAR projection provided as supplementary input. The network estimates a distribution over depth flow field to model the 2D–3D correspondences between the camera and LiDAR data. Finally, the extrinsics are estimated using a PnP solver.
3.1 Problem Formulation
@@
Given the misaligned camera image $I\in\mathbb{R}^{3\times W\times H}$ and LiDAR point cloud $P_{L}\in\mathbb{R}^{3\times N}$, the extrinsic calibration aims to estimate a 6-DoF extrinsic parameter $\mathbf{T}_{est}=\begin{bmatrix}\mathbf{R}_{est}&\mathbf{t}_{est}\\ \mathbf{0}^{T}&1\end{bmatrix}\in SE(3)$ that adjusts the initial extrinsics $\mathbf{T}_{init}\in SE(3)$ for accurate sensor alignment. Here, $W$ and $H$ represent the image width and height, respectively. $\mathbf{R}_{est}\in SO(3)$ denotes the rotation matrix, and $\mathbf{t}_{est}\in\mathbb{R}^{3\times 1}$ represents the translation vector. For matching-based methods, the 2D-3D correspondences between image pixels and LiDAR points are first established. The extrinsic parameter is then estimated via the PnP solver, with RANSAC to prune outliers and enhance robustness.
@@
Figure 3: Comparison of different depth maps. After completion, the depth maps generated by the LiDAR and the camera exhibit similar textures.
3.2 Depth Generation
The main challenge in LiDAR-camera extrinsic calibration lies in cross-modal data matching due to structure and information density differences between point clouds and images. To address this, we propose a novel method that transforms both modalities into dense depth maps, aligning their structure and information density for effective matching. First, we employ MoGe [28] to estimate the dense depth map $D_{I}^{d}\in\mathbb{R}^{1\times W\times H}$ from a camera image. Meanwhile, leveraging the camera’s intrinsic parameter $\mathbf{K}\in\mathbb{R}^{3\times 3}$ and initial extrinsic parameter $\mathbf{T}_{init}=\begin{bmatrix}\mathbf{R}_{init}&\mathbf{t}_{init}\\ \mathbf{0}^{T}&1\end{bmatrix}\in SE(3)$, the LiDAR point clouds $P_{L}$ are projected onto the pixel coordinate system, resulting in a sparse depth map $D_{L}^{s}\in\mathbb{R}^{1\times W\times H}$. The projection process can be formulated as
[…]
where $d_{i}$ represents the depth of each pixel, and $x_{i},y_{i},z_{i},1]^{\mathrm{T}}$ is the homogeneous coordinate of the 3D point, while $u_{i},v_{i},1]^{\mathrm{T}}$ is the corresponding pixel’s homogeneous coordinate in 2D. However, due to the inherent sparsity of LiDAR data, $D_{L}^{s}$ contains only a limited number of valid pixels, leading to a significant information density disparity between $D_{L}^{s}$ and $D_{I}^{d}$, which adversely affects subsequent depth flow estimation. To mitigate this issue, we employ an efficient depth completion method [11] to transform $D_{L}^{s}$ into a dense depth map $D_{L}^{d}\in\mathbb{R}^{1\times W\times H}$. As shown in Fig. 3, this process ensures that both the LiDAR point cloud and the camera image are effectively converted into dense depth maps, facilitating the subsequent depth flow estimation.